# Testing the hypothesis of vector X17 boson by $D$ meson, Charmonium, and $\phi$ meson decays

### v2

@@ Abstract
The recent ATOMKI experiments provided evidence pointing towards the existence of an X17 boson in the anomalous nuclear transitions of Beryllium-8, Helium-4, and Carbon-12. In this work, we consider X17 boson contributions to the previously measured $D$ meson decays which include $D_{s}^{*+}\rightarrow D_{s}^{+}e^{+}e^{-}$, $D_{s}^{*+}\rightarrow D_{s}^{+}\gamma$, $D^{*0}\rightarrow D^{0}e^{+}e^{-}$, and $D^{*0}\rightarrow D^{0}\gamma$, as well as the measured decays of $\psi(2S)\rightarrow\eta_{c}e^{+}e^{-}$, $\psi(2S)\rightarrow\eta_{c}\gamma$, $\phi\rightarrow\eta e^{+}e^{-}$, and $\phi\rightarrow\eta\gamma$. Using the data of the above meson decays, we perform a fitting to the coupling parameters $\varepsilon_{u},\varepsilon_{c}$, and $\varepsilon_{s}$ by treating the couplings $\varepsilon_{u}$ and $\varepsilon_{c}$ as independent from each other rather than assuming the generation universality $\varepsilon_{u}=\varepsilon_{c}$. It is found that the above fitting renders $|\varepsilon_{c}|=6.4\times 10^{-3}$, $|\varepsilon_{s}|=2.0\times 10^{-3}$ and a huge magnitude for $\varepsilon_{u}$, which is in serious tension with $\left|\varepsilon_{u}\right|$ determined from ATOMKI measurements.
@@ I Introduction
In the search for new physics beyond the standard model, the study of low energy processes is an effective avenue besides those focusing on either high energy or precision frontiers. There has been a persistent anomaly in the measured angular distribution of $e^{+}e^{-}$ pairs in the 18.15 MeV decay of the excited state 8Be* by the ATOMKI collaboration over the last few years Krasznahorkay:2015iga ; Krasznahorkay:2017gwn ; Krasznahorkay:2017qfd ; Krasznahorkay:2019lgi . The simplest way to account for the data is to introduce a new X17 boson with a mass in the range of approximately 16 to 17 MeV. The de-excitation of ${}^{8}{\rm Be}^{*}$ is assumed to emit a X17 boson which then decays to the $e^{+}e^{-}$ pair as detected by the above experiments. It is desirable to confirm the existence of X17 boson which undoubtedly leads to new physics beyond the Standard Model (SM).
@@
where $q$ stands for $u$, $d$ and $s$ quarks, while $Q$ represents heavy quarks. From the data of ${}^{8}{\rm Be}$ anomaly, $R_{X}=5.8\times 10^{-6}$, with the assumption ${\rm BR}\left(X\rightarrow e^{+}e^{-}\right)=1$, it has been established that Feng:2016jff ; Fornal:2017msy $\varepsilon_{u}\simeq\pm 3.7\times 10^{-3}$ and $\varepsilon_{d}\simeq\mp 7.4\times 10^{-3}$, where the null result of $\pi^{0}\to X\gamma$ search by NA48/2 NA482:2015wmo has been used. Further refined analysis takes into account isospin effects and their mixings Feng:2016ysn . Such an analysis considers three distinct isospin scenarios: no isospin effects, isospin mixing, and isospin mixing & breaking. Assuming that X17 is a vector boson, a statistical $\chi^{2}$ test based upon data from anomalous transitions observed in excited ${}^{8}\mathrm{Be},{}^{4}\mathrm{He}$, and ${}^{12}\mathrm{C}$ has been performed in Denton:2023gat . Utilizing eleven measurements that include angular and width data from the anomalous transitions in ATOMKI experiments, and the null result of $\pi^{0}\to X\gamma$ search NA482:2015wmo under the assumption ${\rm BR}\left(X\rightarrow e^{+}e^{-}\right)=1$, the above $\chi^{2}$ test determines the ranges of three parameters, $m_{X},\varepsilon_{n},\,{\rm and}\,\varepsilon_{p}$ where the latter two are related to the quark-level couplings by $\varepsilon_{p}=2\varepsilon_{u}+\varepsilon_{d}$ and $\varepsilon_{n}=2\varepsilon_{d}+\varepsilon_{u}$. It was found that  Denton:2023gat $\left|\varepsilon_{u}\right|\simeq(0.5-0.9)\times 10^{-3}$ and $\left|\varepsilon_{d}\right|\simeq(2.5-2.9)\times 10^{-3}$ with $\varepsilon_{u}$ and $\varepsilon_{d}$ in opposite signs. Specifically the most favored parameter set is $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$ and $m_{X}=16.84\ {\rm MeV}$ if isospin effects are not considered. With both isospin mixing and breaking effects considered, the most favored parameter set becomes $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$ and $m_{X}=16.85\ {\rm MeV}$. It is noteworthy that the above fitted values for $\left|\varepsilon_{u}\right|$ and $\left|\varepsilon_{d}\right|$ are much smaller than those determined earlier.
@@
Given the hypothesis of X17 boson to account for the anomalous decays of ${}^{8}\mathrm{Be},{}^{4}\mathrm{He}$, and ${}^{12}\mathrm{C}$ excited states, it is of interest to test this hypothesis on heavy hadrons which involve the couplings of X17 boson to the second and the third generation of quarks, in addition to the couplings $\varepsilon_{u}$ and $\varepsilon_{d}$. In fact, ${}^{8}{\rm Be}$ anomaly already motivated the study of X17 boson effect to $K^{+}\to\mu^{+}\nu_{\mu}e^{+}e^{-}$ decay through the interaction/decay chain, $K^{+}\to\mu^{+}\nu_{\mu}X\to\mu^{+}\nu_{\mu}e^{+}e^{-}$ Chiang:2016cyf with the assumption $\varepsilon_{s}=\varepsilon_{d}$. It was also proposed that X17 boson can be searched by BESIII or Belle II by the decay chain $J/\Psi\to\eta_{c}X^{*}\to\eta_{c}e^{+}e^{-}$ Ban:2020uii , which is an excess to the regular decay chain $J/\Psi\to\eta_{c}\gamma^{*}\to\eta_{c}e^{+}e^{-}$. Furthermore, effects of X17 boson to decays of $D^{*(+,0)}$, $D^{*+}_{s}$, and the corresponding $B$ mesons were also studied in Castro:2021gdf . In the latter work, it was stated that the theoretical prediction of $R_{D_{s}^{*+}}\equiv\Gamma(D_{s}^{*+}\rightarrow D^{+}_{s}e^{+}e^{-})/\Gamma(D_{s}^{*+}\rightarrow D^{+}_{s}\gamma)$ with the contribution $D_{s}^{*+}\to D^{+}_{s}+X\to D^{+}_{s}+e^{+}+e^{-}$ included in the numerator is consistent with the CLEO measurement $R_{D_{s}^{*+}}=\left(7.2_{-1.6}^{+1.8}\right)\times 10^{-3}$ CLEO:2011mla . We note that the authors of Ref. Castro:2021gdf calculated the X17 contribution to $R_{D_{s}^{*+}}$ by taking $\varepsilon_{u}=\pm 3.7\times 10^{-3}$ and $\varepsilon_{d}=\mp 7.4\times 10^{-3}$  Feng:2016jff and assuming $\varepsilon_{s}=\varepsilon_{d}$, and $\varepsilon_{c}=\varepsilon_{u}$. Furthermore the signals due to X17 boson in other decay modes of $D^{*}$ and $B^{*}$ mesons were predicted.
@@
In this article, we shall revisit the contributions of X17 boson to $D^{*0}$ and $D^{*+}_{s}$ meson decays since there existed a measurement on $R_{D^{*0}}\equiv\Gamma(D^{*0}\rightarrow D^{0}e^{+}e^{-})/\Gamma(D^{*0}\rightarrow D^{0}\gamma)$ after the publication of Castro:2021gdf . The measurement gives $R_{D^{*0}}=(11.08\pm 0.9)\times 10^{-3}$ BESIII:2021vyq . Furthermore, we are motivated to tackle this issue because the updated values for $\left|\varepsilon_{u}\right|$ and $\left|\varepsilon_{d}\right|$ Denton:2023gat are much smaller than those determined earlier Feng:2016jff . We found that the calculation on $R_{D_{s}^{*+}}$ by Castro:2021gdf is incorrect. After correcting the calculation, we found that the prediction of $R_{D_{s}^{*+}}$ is consistent with CLEO data only for the updated values of $\varepsilon_{c}\equiv\varepsilon_{u}$ and $\varepsilon_{s}\equiv\varepsilon_{d}$. On the other hand, the data on $R_{D^{*0}}$ requires a much larger $\varepsilon_{u}$ with the assumption $\varepsilon_{c}=\varepsilon_{u}$. It is obvious that the updated ranges of $\varepsilon_{u}$ and $\varepsilon_{d}$ with the assumptions $\varepsilon_{c}=\varepsilon_{u}$, and $\varepsilon_{s}=\varepsilon_{d}$ cannot simultaneously account for $R_{D_{s}^{*+}}$ and $R_{D^{*0}}$. To ease the tension, we note that the contribution of X17 boson to $R_{D_{s}^{*+}}$ behaves approximately as $\left|\varepsilon_{c}+2\varepsilon_{s}\right|$ with $\varepsilon_{c}$ and $\varepsilon_{s}$ in opposite signs as we shall see in Fig. 3 later. Hence one can account for both $R_{D^{*0}}$ and $R_{D_{s}^{*+}}$ by raising both $\left|\varepsilon_{c}\right|$ and $\left|\varepsilon_{s}\right|$ while maintaining $\left|\varepsilon_{c}+2\varepsilon_{s}\right|$ fixed. On the other hand, this implies that the assumption $\varepsilon_{s}=\varepsilon_{d}$ is no longer appropriate. In view of this, one may lift the restriction $\varepsilon_{c}=\varepsilon_{u}$ as well. Finally, we note that the enhancement of both $\varepsilon_{c}$ and $\varepsilon_{s}$ should impact the decays of charmonium and $\phi$ meson. Therefore we also include $\psi(2S)\rightarrow\eta_{c}e^{+}e^{-}$, $\psi(2S)\rightarrow\eta_{c}\gamma$, $\phi\rightarrow\eta e^{+}e^{-}$, and $\phi\rightarrow\eta\gamma$ in our analysis. We shall perform fittings to the coupling strengths $\varepsilon_{u},\varepsilon_{c}$, and $\varepsilon_{s}$ with the data of $R_{D_{s}^{*+}}$, $R_{D^{*0}}$, $R_{\psi(2S)}\equiv\Gamma(\psi(2S)\rightarrow\eta_{c}e^{+}e^{-})/\Gamma(\psi(2S)\rightarrow\eta_{c}\gamma)$ BESIII:2022jde , and $R_{\phi}\equiv\Gamma(\phi\rightarrow\eta e^{+}e^{-})/\Gamma(\phi\rightarrow\eta\gamma)$ Achasov:2000ne ; CMD-2:2000qla ; KLOE-2:2014hfk . It is important to compare the favored value of $\varepsilon_{u}$ in this fitting to that obtained from the decay anomalies of excited 8Be, 4He, and 12C. Here we fix the mass of X17 boson at $16.85$ MeV which is the best-fit value of $m_{X}$ obtained in Denton:2023gat . We adopt this simplification since various decay rates considered here are rather insensitive to $m_{X}$ over its allowed range.
@@
This article is organized as follows. In Sec. II, we review previous studies on the X17 contributions to heavy meson decays $H^{*}\to He^{+}e^{-}$ normalized by $H^{*}\to H\gamma$ with $H^{*}$ representing $D^{*0}$, $D^{*+}$, and $D^{*+}_{s}$ vector mesons and $H$ their pseudoscalar counterparts. We note that the semi-leptonic decay $H^{*}\to He^{+}e^{-}$ proceeds through both $H^{*}\to H\gamma^{*}\to He^{+}e^{-}$ and $H^{*}\to HX\to He^{+}e^{-}$. The $H^{*}\to H\gamma^{*}$ form factor is essentially $\langle H|J^{\rm em}_{\mu}|H^{*}\rangle$ with $J^{\rm em}_{\mu}=ee_{Q}\bar{Q}\gamma_{\mu}Q+ee_{q}\bar{q}\gamma_{\mu}q$. The evaluation of this matrix element was discussed in Refs. Cho:1992nt ; Amundson:1992yp ; Cheng:1992xi ; Colangelo:1993zq . It is well known that $\langle H|\bar{Q}\gamma_{\mu}Q|H^{*}\rangle$ is fixed by heavy-quark (HQ) symmetry Isgur-Wise while various approaches are employed to compute $\langle H|\bar{q}\gamma_{\mu}q|H^{*}\rangle$. Here we apply the model of vector meson dominance (VMD) BKY proposed in Ref. Colangelo:1993zq for computing the matrix element of light quark current. This choice enables the comparison of our result with that of Ref. Castro:2021gdf which also employs VMD model for computing $\langle H|\bar{q}\gamma_{\mu}q|H^{*}\rangle$. For the $H^{*}\to HX$ transition, the matrix element $\langle H|J^{X}_{\mu}|H^{*}\rangle$ with $J^{X}_{\mu}=e\varepsilon_{Q}\bar{Q}\gamma_{\mu}Q+e\varepsilon_{q}\bar{q}\gamma_{\mu}q$ can be easily inferred from $\langle H|J^{\rm em}_{\mu}|H^{*}\rangle$. We shall review the details of calculating $R_{D_{s}^{*+}}$ and correct the error made in Ref. Castro:2021gdf . In the same section, we shall also discuss the implication of $R_{D^{*0}}$ on values of $\varepsilon_{c}$ and $\varepsilon_{u}$. It will be demonstrated that the conditions $\varepsilon_{c}=\varepsilon_{u}$ and $\varepsilon_{d}=\varepsilon_{s}$ should be relaxed in order that the data of $R_{D_{s}^{*+}}$ and $R_{D^{*0}}$ can be simultaneously accounted for. In Sec. III, we present the fitting to the couplings $\varepsilon_{u}$, $\varepsilon_{s}$, and $\varepsilon_{c}$ with the data of $\psi(2S)\rightarrow\eta_{c}e^{+}e^{-}$, $\psi(2S)\rightarrow\eta_{c}\gamma$, $\phi\rightarrow\eta e^{+}e^{-}$, $\phi\rightarrow\eta\gamma$, and that of $R_{D_{s}^{*}}$ and $R_{D^{*0}}$. Sec. IV addresses the implications of our fitting and the outlook.

### v3

@@ Abstract
The recent ATOMKI experiments provided evidence pointing towards the existence of an X17 boson in the anomalous nuclear transitions of Beryllium-8, Helium-4, and Carbon-12. In this work, we consider X17 boson contributions to the previously measured $D$ meson decays which include $D_{s}^{*+}\rightarrow D_{s}^{+}e^{+}e^{-}$, $D_{s}^{*+}\rightarrow D_{s}^{+}\gamma$, $D^{*0}\rightarrow D^{0}e^{+}e^{-}$, and $D^{*0}\rightarrow D^{0}\gamma$, as well as the measured decays of $\psi(2S)\rightarrow\eta_{c}e^{+}e^{-}$, $\psi(2S)\rightarrow\eta_{c}\gamma$, $\phi\rightarrow\eta e^{+}e^{-}$, and $\phi\rightarrow\eta\gamma$. Using the data of the above meson decays, we perform a fitting to the coupling parameters $\varepsilon_{u},\varepsilon_{c}$, and $\varepsilon_{s}$ by treating the couplings $\varepsilon_{u}$ and $\varepsilon_{c}$ as independent from each other rather than assuming the generation universality $\varepsilon_{u}=\varepsilon_{c}$. It is found that the above fitting renders $|\varepsilon_{c}|=7.6\times 10^{-3}$, $|\varepsilon_{s}|=2.4\times 10^{-3}$ and a huge magnitude for $\varepsilon_{u}$, which is in serious tension with $\left|\varepsilon_{u}\right|$ determined from ATOMKI measurements. Using our fitted range for $\varepsilon_{c}$ and the range for $\varepsilon_{d}$ from ATOMKI measurements, we predict the range for $D^{*+}\rightarrow D^{+}e^{+}e^{-}$ decay rate.
@@ I Introduction
In the search for new physics beyond the Standard Model (SM), the study of low energy processes is an effective avenue besides those focusing on either high energy or precision frontiers. There has been a persistent anomaly in the measured angular distribution of $e^{+}e^{-}$ pairs in the 18.15 MeV decay of the excited state 8Be* by the ATOMKI collaboration over the last few years Krasznahorkay:2015iga ; Krasznahorkay:2017gwn ; Krasznahorkay:2017qfd ; Krasznahorkay:2019lgi . The simplest way to account for the data is to introduce a new X17 boson with a mass in the range of approximately 16 to 17 MeV. The de-excitation of $^{8}{\rm Be}^{*}$ is assumed to emit a X17 boson which then decays to the $e^{+}e^{-}$ pair as detected by the above experiments. It is desirable to confirm the existence of X17 boson which undoubtedly leads to new physics beyond SM.
@@
where $q$ stands for $u$, $d$ and $s$ quarks, while $Q$ represents heavy quarks. From the data of $^{8}{\rm Be}$ anomaly, $R_{X}=5.8\times 10^{-6}$, with the assumption ${\rm BR}\left(X\rightarrow e^{+}e^{-}\right)=1$, it has been established that Feng:2016jff ; Fornal:2017msy $\varepsilon_{u}\simeq\pm 3.7\times 10^{-3}$ and $\varepsilon_{d}\simeq\mp 7.4\times 10^{-3}$, where the null result of $\pi^{0}\to X\gamma$ search by NA48/2 NA482:2015wmo has been used. Further refined analysis takes into account isospin effects and their mixings Feng:2016ysn . Such an analysis considers three distinct isospin scenarios: no isospin effects, isospin mixing, and isospin mixing & breaking. Assuming that X17 is a vector boson, a statistical $\chi^{2}$ test based upon data from anomalous transitions observed in excited $^{8}\mathrm{Be},{}^{4}\mathrm{He}$, and $^{12}\mathrm{C}$ has been performed in Denton:2023gat . Utilizing eleven measurements that include angular and width data from the anomalous transitions in ATOMKI experiments, and the null result of $\pi^{0}\to X\gamma$ search NA482:2015wmo under the assumption ${\rm BR}\left(X\rightarrow e^{+}e^{-}\right)=1$, the above $\chi^{2}$ test determines the ranges of three parameters, $m_{X},\varepsilon_{n},\,{\rm and}\,\varepsilon_{p}$ where the latter two are related to the quark-level couplings by $\varepsilon_{p}=2\varepsilon_{u}+\varepsilon_{d}$ and $\varepsilon_{n}=2\varepsilon_{d}+\varepsilon_{u}$. It was found that  Denton:2023gat $\left|\varepsilon_{u}\right|\simeq(0.5-0.9)\times 10^{-3}$ and $\left|\varepsilon_{d}\right|\simeq(2.5-2.9)\times 10^{-3}$ with $\varepsilon_{u}$ and $\varepsilon_{d}$ in opposite signs. Specifically, the most favored parameter set is $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$ and $m_{X}=16.84\ {\rm MeV}$ if isospin effects are not considered. With both isospin mixing and breaking effects considered, the most favored parameter set becomes $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$ and $m_{X}=16.85\ {\rm MeV}$. It is noteworthy that the above fitted values for $\left|\varepsilon_{u}\right|$ and $\left|\varepsilon_{d}\right|$ are much smaller than those determined earlier.
@@
Given the hypothesis of X17 boson to account for the anomalous decays of $^{8}\mathrm{Be},{}^{4}\mathrm{He}$, and $^{12}\mathrm{C}$ excited states, it is of interest to test this hypothesis on heavy hadrons which involve the couplings of X17 boson to the second and the third generation of quarks, in addition to the couplings $\varepsilon_{u}$ and $\varepsilon_{d}$. In fact, $^{8}{\rm Be}$ anomaly already motivated the study of X17 boson effect to $K^{+}\to\mu^{+}\nu_{\mu}e^{+}e^{-}$ decay through the interaction/decay chain, $K^{+}\to\mu^{+}\nu_{\mu}X\to\mu^{+}\nu_{\mu}e^{+}e^{-}$ Chiang:2016cyf with the assumption $\varepsilon_{s}=\varepsilon_{d}$. It was also proposed that X17 boson can be searched by BESIII or Belle II by the decay chain $J/\Psi\to\eta_{c}X^{*}\to\eta_{c}e^{+}e^{-}$ Ban:2020uii , which is an excess to the regular decay chain $J/\Psi\to\eta_{c}\gamma^{*}\to\eta_{c}e^{+}e^{-}$. Furthermore, effects of X17 boson to decays of $D^{*(+,0)}$, $D^{*+}_{s}$, and the corresponding $B$ mesons were also studied in Castro:2021gdf . In the latter work, it was stated that the theoretical prediction of $R_{D_{s}^{*+}}\equiv\Gamma(D_{s}^{*+}\rightarrow D^{+}_{s}e^{+}e^{-})/\Gamma(D_{s}^{*+}\rightarrow D^{+}_{s}\gamma)$ with the contribution $D_{s}^{*+}\to D^{+}_{s}+X\to D^{+}_{s}+e^{+}+e^{-}$ included in the numerator is consistent with the CLEO measurement $R_{D_{s}^{*+}}=\left(7.2_{-1.6}^{+1.8}\right)\times 10^{-3}$ CLEO:2011mla . We note that the authors of Ref. Castro:2021gdf calculated the X17 contribution to $R_{D_{s}^{*+}}$ by taking $\varepsilon_{u}=\pm 3.7\times 10^{-3}$ and $\varepsilon_{d}=\mp 7.4\times 10^{-3}$  Feng:2016jff and assuming $\varepsilon_{s}=\varepsilon_{d}$, and $\varepsilon_{c}=\varepsilon_{u}$. Furthermore, the signals due to X17 boson in other decay modes of $D^{*}$ and $B^{*}$ mesons were predicted.
@@
In this article, we shall revisit the contributions of X17 boson to $D^{*0}$ and $D^{*+}_{s}$ meson decays since there existed a measurement on $R_{D^{*0}}\equiv\Gamma(D^{*0}\rightarrow D^{0}e^{+}e^{-})/\Gamma(D^{*0}\rightarrow D^{0}\gamma)$ after the publication of Castro:2021gdf . The measurement gives $R_{D^{*0}}=(11.08\pm 0.90)\times 10^{-3}$ BESIII:2021vyq . Furthermore, we are motivated to tackle this issue because the updated values for $\left|\varepsilon_{u}\right|$ and $\left|\varepsilon_{d}\right|$ Denton:2023gat are much smaller than those determined earlier Feng:2016jff . We found that the calculation on $R_{D_{s}^{*+}}$ by Castro:2021gdf is incorrect. After correcting the calculation, we found that the prediction of $R_{D_{s}^{*+}}$ is consistent with CLEO data only for the updated values of $\varepsilon_{c}\equiv\varepsilon_{u}$ and $\varepsilon_{s}\equiv\varepsilon_{d}$. On the other hand, the data on $R_{D^{*0}}$ requires a much larger $\varepsilon_{u}$ with the assumption $\varepsilon_{c}=\varepsilon_{u}$. It is obvious that the updated ranges of $\varepsilon_{u}$ and $\varepsilon_{d}$ with the assumptions $\varepsilon_{c}=\varepsilon_{u}$, and $\varepsilon_{s}=\varepsilon_{d}$ cannot simultaneously account for $R_{D_{s}^{*+}}$ and $R_{D^{*0}}$. To ease the tension, we note that the contribution of X17 boson to $R_{D_{s}^{*+}}$ behaves approximately as $\left|\varepsilon_{c}+2\varepsilon_{s}\right|$ with $\varepsilon_{c}$ and $\varepsilon_{s}$ in opposite signs as we shall see in Fig. 3 later. Hence, one can account for both $R_{D^{*0}}$ and $R_{D_{s}^{*+}}$ by raising both $\left|\varepsilon_{c}\right|$ and $\left|\varepsilon_{s}\right|$ while maintaining $\left|\varepsilon_{c}+2\varepsilon_{s}\right|$ fixed. On the other hand, this implies that the assumption $\varepsilon_{s}=\varepsilon_{d}$ is no longer appropriate. In view of this, one may lift the restriction $\varepsilon_{c}=\varepsilon_{u}$ as well. Finally, we note that the enhancement of both $\varepsilon_{c}$ and $\varepsilon_{s}$ should impact the decays of charmonium and $\phi$ meson. Therefore, we also include $\psi(2S)\rightarrow\eta_{c}e^{+}e^{-}$, $\psi(2S)\rightarrow\eta_{c}\gamma$, $\phi\rightarrow\eta e^{+}e^{-}$, and $\phi\rightarrow\eta\gamma$ in our analysis. We shall perform fittings to the coupling strengths $\varepsilon_{u},\varepsilon_{c}$, and $\varepsilon_{s}$ with the data of $R_{D_{s}^{*+}}$, $R_{D^{*0}}$, $R_{\psi(2S)}\equiv\Gamma(\psi(2S)\rightarrow\eta_{c}e^{+}e^{-})/\Gamma(\psi(2S)\rightarrow\eta_{c}\gamma)$ BESIII:2022jde , and $R_{\phi}\equiv\Gamma(\phi\rightarrow\eta e^{+}e^{-})/\Gamma(\phi\rightarrow\eta\gamma)$ Achasov:2000ne ; CMD-2:2000qla ; KLOE-2:2014hfk . It is important to compare the favored value of $\varepsilon_{u}$ in this fitting to that obtained from the decay anomalies of excited 8Be, 4He, and 12C. Here, we fix the mass of X17 boson at $16.85$ MeV which is the best-fit value of $m_{X}$ obtained in Denton:2023gat . We adopt this simplification since various decay rates considered here are rather insensitive to $m_{X}$ over its allowed range.
@@
This article is organized as follows. In Sec. II, we review previous studies on the X17 contributions to heavy meson decays $H^{*}\to He^{+}e^{-}$ normalized by $H^{*}\to H\gamma$ with $H^{*}$ representing $^{*0}, D+D^{*+}$, and $D^{*+}_{s}$ vector mesons and $H$ their pseudoscalar counterparts. We note that the semi-leptonic decay $H^{*}\to He^{+}e^{-}$ proceeds through both $H^{*}\to H\gamma^{*}\to He^{+}e^{-}$ and $H^{*}\to HX\to He^{+}e^{-}$. The $H^{*}\to H\gamma^{*}$ form factor is essentially $\langle H|J^{\rm em}_{\mu}|H^{*}\rangle$ with $J^{\rm em}_{\mu}=ee_{Q}\bar{Q}\gamma_{\mu}Q+ee_{q}\bar{q}\gamma_{\mu}q$. The evaluation of this matrix element was discussed in Refs. Cho:1992nt ; Amundson:1992yp ; Cheng:1992xi ; Colangelo:1993zq . It is well known that $\langle H|\bar{Q}\gamma_{\mu}Q|H^{*}\rangle$ is fixed by heavy-quark (HQ) symmetry Isgur-Wise while various approaches are employed to compute $\langle H|\bar{q}\gamma_{\mu}q|H^{*}\rangle$. Here, we apply the model of vector meson dominance (VMD) BKY proposed in Ref. Colangelo:1993zq for computing the matrix element of light quark current. This choice enables the comparison of our result with that of Ref. Castro:2021gdf which also employs VMD model for computing $\langle H|\bar{q}\gamma_{\mu}q|H^{*}\rangle$. For the $H^{*}\to HX$ transition, the matrix element $\langle H|J^{X}_{\mu}|H^{*}\rangle$ with $J^{X}_{\mu}=e\varepsilon_{Q}\bar{Q}\gamma_{\mu}Q+e\varepsilon_{q}\bar{q}\gamma_{\mu}q$ can be easily inferred from $\langle H|J^{\rm em}_{\mu}|H^{*}\rangle$. We shall review the details of calculating $R_{D_{s}^{*+}}$ and correct the error made in Ref. Castro:2021gdf . In the same section, we shall also discuss the implication of $R_{D^{*0}}$ on values of $\varepsilon_{c}$ and $\varepsilon_{u}$. It will be demonstrated that the conditions $\varepsilon_{c}=\varepsilon_{u}$ and $\varepsilon_{d}=\varepsilon_{s}$ should be relaxed in order that the data of $R_{D_{s}^{*+}}$ and $R_{D^{*0}}$ can be simultaneously accounted for. In Sec. III, we present the fitting to the couplings $\varepsilon_{u}$, $\varepsilon_{s}$, and $\varepsilon_{c}$ with the data of $\psi(2S)\rightarrow\eta_{c}e^{+}e^{-}$, $\psi(2S)\rightarrow\eta_{c}\gamma$, $\phi\rightarrow\eta e^{+}e^{-}$, $\phi\rightarrow\eta\gamma$, and that of $R_{D_{s}^{*}}$ and $R_{D^{*0}}$. We also predict the range for $R_{D^{*+}}$ using the $\varepsilon_{c}$ range given by the above fitting and the range for $\varepsilon_{d}$ given by Ref. Denton:2023gat . Sec. IV addresses the implications of our results and the outlook.